Start learning →

If $P(E) = 0.35$, then $P(\text{not } E)$ is

A0.35
B1.35
C0.65
D0.35 again
Answer & Solution
Correct answer: C. 0.65
1. Use the complement rule: $P(\text{not }E) = 1 - P(E)$. 2. $= 1 - 0.35 = 0.65$. 3. So $P(\text{not }E) = 0.65$. _Source: Karnataka SSLC (KSEEB) Class 10 Mathematics, Ch14 'Probability', §14.1_
Practise this question with Quill — Karnataka SSLC (Class 10) →

Keep practising Probability

Browse all Probability questions for Karnataka SSLC (Class 10), or explore more Mathematics topics.

Start a Probability lesson →
Related questions